derivation for body projected vertically upwards from the top of the tower, a) maximum height, B)time spent by the body above the top the tower, c)body reaches the ground with avelocity d) maximum height above the ground

Grade:10

1 Answers

Vikas TU
6 years ago
Dear Student,
Derivationw ould go like this:
Tower ht=h1
v=u+at
=>0=u-gt
=>t=u/g
max ht=ut-gt2/2
=>h=u2/2g
time to go up=u/g
time to come back=u/g
total time=2u/g
v2=u2+2gh1
=> velocity with which the body reaches the ground= v = √(u^2+2gh1)
max ht above the ground = h+h1
=u2/2g + h1.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)